OpenStudy (anonymous):

Suppose that 67% of the students at a college are female and 50% are white. Therefore, if F is the event that a randomly chosen student is female, and W is the event that a randomly chosen student is white, then Pr(F) = 0.67 and Pr(W) = 0.5. Assuming that F and W are not necessarily independent events, what is the maximum percentage of the student body that could be both white and female?

OpenStudy (anonymous):

Since there is no information given that says event W can't be completely inside of event F, you could assume that there is a chance all of the white students are also female. Therefore, the maximum percentage P(F and W) = 0.50

OpenStudy (mathmate):

\(P(F\cap W)=P(F)+P(W)-P(F\cup W)\) \(P(F\cap W)=0.67+0.5-P(F\cup W)\) For \(P(F\cap W)\) to be a maximum, \(P(F\cup W)\) must be a minimum, or 0.67. Therefore \(P(F\cap W)=0.67+0.5-0.67=0.5\) exactly as @derrowap had it.

OpenStudy (anonymous):

@mathmate First line: Why does F and W = F+W-(F or W) -- doesn't that just leave F or W?

OpenStudy (mathmate):

 The original equation is F or W = F + W - F and W because F and W is counted twice in F+W. We transpose it to get what I posted. I hope that explains it. See also the drawing above.